I know the solution behind this question but didn't understand how it is calculated and by what logic. If the swimmer is swimming at normal to the river, then shouldn't it's path be the hypotension. Why is it taken as 1km? I need to understand in depth. 
Ncert question: a man can swim with a speed of 4 km/hr in still water. How long does he take to cross a river 1km wide if the river flows steadily at 3km/hr and he makes his strokes normal to the river current? How far down the river does he go when he reaches the other bank? 

Grade:11

1 Answers

Arun
4 years ago
The man goes normal to the flow of the current of the river. So , the net velocity of the man will be the resultant of the man's velocity and the perpendicular velocity of the current. Now the man has to cover the width of the river that is 1 km, so the time taken by the man will be,
t = 1/4 = 0.25 seconds.
 
Vertical distance travelled due to the current = 3*0.25 = 0.75 km
so the man lands at a distance of 0.75 km below from the river.